# ETH-monotonicity and the black hole singularity

### 11 Mar 2026 v3

@@ 3 ETH-monotonicity
In ETHmono ; sorokhaibam2025ethmono2D , we found some interesting constraints on $\bar{E}$-dependence of the $f$-function in (1.1). These constraints are collectively called ETH-monotonicity. These constraints arise from a non-trivial relation between Berry’s conjecture and the second law of thermodynamics in the form of the Kelvin statement. It implies that, when the system is perturbed starting from an energy eigenstate or a microcanonical ensemble, more energy is absorbed than an equivalent canonical ensemble. The systems considered were lattice models with various energy scales. So, we will redefine ETH-monotonicity for holographic CFTs below. The main criteria is that an energy eigenstate absorbs more energy than an equivalent canonical ensemble, when the system is perturbed. The low energy eigenstates are of particular interest. The high energy eigenstates gain energy same as their canonical ensemble counterpart due to equivalence of ensembles. Even though we got the inspiration from the above previous works, the present work is largely independent.
We define ETH-monotonicity for holographic CFTs as follows:
@@
(3.2) is the most severe constraint, $\eta=1$ is necessary. We emphasize that while taking $\bar{E}\to 0$ we are not shrinking the horizon radius to the string length scale $l_{s}$ or the Planck length scale $l_{P}$. We will briefly comment on the quantum gravity limit in the next section 4.
We assume $\bar{E}=0$ is the ground state because the Casimir energy does not participate in the dynamical set-up that we are considering. All CFTs do not have any intrinsic system size scale. So, R3 is formulated in terms of $\bar{E}$. R3 was originally formulated in terms of the degrees of freedom $L^{d}$, where $L$ is the system size and $d$ is the spatial dimension of the system.
The perturbation is by a source term $\lambda(t)\mathcal{O}$. Starting from a canonical ensemble with energy $\bar{E}_{\beta}$, the gain in energy by the system at leading order in perturbative expansion is ETHmono
@@
$\tilde{\lambda}(\omega)$ is the Fourier transform of $\lambda(t)$. For a hermitian operator, $\lambda(t)$ is real and $\tilde{\lambda}(\omega)=\tilde{\lambda}(-\omega)^{*}$. The above quantity is identically positive (negative or zero) for positive (negative or infinite) temperature. So, the second law of thermodynamics is guaranteed in canonical ensembles (in general, for any passive density matrix) Polkovnikov_2008 . The integral is finite because the $f$-function falls at least as fast as $e^{-|\beta\omega|/4}$, otherwise the theory is not UV complete. $\Delta E_{\beta}$ is of order $O(1)$ because we are perturbing the system with a light operator.
@@
We have used Taylor expansions of $S(E_{n}+\omega)$ and $S(E_{n}+\omega/2)$ about the energy of the initial state $E_{n}$, $\omega=E_{m}-E_{n}$, and $\beta=S^{\prime}(\bar{E})|_{\bar{E}=E_{n}}$ is the effective inverse temperature at $E_{n}$. The difference with the expression of $\Delta E_{\beta}$ is in the $\bar{E}$-dependence. For a finite system (not thermodynamic limit), $f(\bar{E})$ is an increasing function of $\bar{E}$. So, the system gains more energy if the initial state is an eigenstate, over and above the entropic contribution. We have neglected the contributions of the higher derivative terms $S^{\prime\prime}(\bar{E})$, $S^{\prime\prime\prime}(\bar{E})$, etc. They are suppressed by increasing positive powers of $G_{N}$ factor. But $S^{\prime\prime}(\bar{E})$ can compete with ETH-monotonicity and we will study this quantum fluctuations term in details in section 5.
(3.7) also holds true in a microcanonical ensemble with energy expectation value $\bar{E}_{mc}$, see Appendix A for details. The microcanonical energy window $\delta E$ is of the order of $O(1/\sqrt{G_{N}})$. But $\delta E$ is usually taken to be parametrically much smaller than $\sqrt{G_{N}}\bar{E}_{mc}\sim 1/\sqrt{G_{N}}$ Srednicki_1999 . The correction from the microcanonical energy window $\delta E$ arises from two different terms - one from the entropy exponentials and the other from $f(\bar{E})$. Both terms are of the order of $O(G^{3/2}_{N})$ so $\delta E$ which measures energy fluctuation in microcanonical ensemble does not affect ETH-monotonicity in any dimension. This is expected because ETH-monotonicity arises due to inequivalence of ensembles and we are working with microcanonical ensemble.
In holographic CFTs, $\bar{E}$ is order $1/\tilde{G}_{N}$ times larger than $\omega$. So, we can Taylor expand $f(E_{n}+\omega/2,\omega)$ about $(E_{n},\omega)$ in (3.7). Using the Taylor expansion, we can extract the extra contribution to the gain in energy from ETH-monotonicity when the system is initially in a microcanonical ensemble. Let us denote that quantity by $\Delta E_{mc,f}$ and it is given by
@@
The extra gain in energy is $O(G^{1}_{N})$ because it is controlled by $f^{\prime}(\bar{E})/f(\bar{E})$, compared to $\Delta E_{\beta}$. So $f^{\prime}(\bar{E})/f(\bar{E})$ is the main object of our interest.
Pure states: In place of the microcanonical ensemble, one can also consider pure states $|\Psi\rangle=\sum_{n}c_{n}|n\rangle$ with $|n\rangle$ within the energy window $\delta E$ of the order $O(1/\sqrt{G_{N}})$. These states have been called typical states in Srednicki_1999 . For the change in energy, the calculation is same as in the microcanonical ensemble because the cross terms $\mathcal{O}_{n_{1}m}\mathcal{O}_{mn_{2}}$ with $n_{1}\neq n_{2}$ in (3.5) drops out due to the pseudo random factor $R_{n_{1}m}R_{mn_{2}}$ averaging out to zero.
The motivations for ETH-monotonicity are as follows. R1 follows from equivalence of ensembles in the themodynamic limit $\bar{E}\to\infty$. (3.3) of R3 is the mathematical expression. It implies $\Delta E_{mc}\to\Delta E_{\beta}$ in the thermodynamic limit. R2 is generally desirable because it will make an eigenstate gain more energy, not less, than from pure entropic consideration. Moreover, many observables in quantum chaotic systems are found to obey R2 in ETHmono ; sorokhaibam2025ethmono2D .
We have special interest in the limit $\bar{E}\to 0$ because it corresponds to small black holes with high curvature at the horizon, and ultimately the black hole singularity. Naively, we expect that the extra gain in energy would persist in the limit $\bar{E}\to 0$ if $\eta>0$ in (3.2). For example, we will find that 2d2-d holographic CFT has $\eta=3/2,\kappa=\pi\omega/4$. But with this value of $\eta$, $f(\bar{E})$ is
@@
where $f_{0}>0$ is the integration constant. The problem with this function is that $\Delta E_{mc}$ or $\Delta E_{\beta}$ vanishes exponentially fast. There could be power-law dependent factor of $\bar{E}$ but it is not important for $f^{\prime}(\bar{E})/f(\bar{E})$ in the presence of the above exponential. On the other hand, if $\eta=1$, then $f(\bar{E})$ is
$f(\bar{E})=f_{0}\bar{E}^{\kappa}$ (3.11)
If $\eta$ is less than $1$, say $\eta=1/2$, then $f(\bar{E})=f_{0}e^{2\kappa\sqrt{\tilde{G}_{N}\bar{E}}}$ which is not possible because $\bar{E}=0$ is the ground state. $f(\bar{E})$ signifies the strength of quantum fluctuations mimicking thermal fluctuations Srednicki_1999 and it is expected to be equal to zero at zero temperature. So in conclusion, $\eta=1$ is a very strict constraint in (3.2).
Note that $\eta>0$, in particular $\eta=1$, implies that ETH-monotonicity can start competing with the entropic factor in the limit $\bar{E}\to 0$. We can examine this competition by calculating the relative extra gain in energy at a particular frequency $\omega$ defined by
@@ 4 Summary of results
Our technical results are:
@@
    In Higher-d holographic CFTs, the relative extra gain in energy $\overline{\Delta E}_{n}$ measures the curvature at the black hole horizon of small black holes,
    $\lim_{\bar{E}\to 0}\overline{\Delta E}_{mc}=\frac{\kappa(d-2)\,\tilde{G}_{N}}{2\pi r_{h}^{d-1}}.$ (4.2)
    The $\omega$-dependence drops out. Note that $\overline{\Delta E}_{n}$ is a natural quantity which one can even measure in a laboratory for any chaotic system. ETH-monotonicity also dominates the quantum fluctuations in the small black hole microstates. ETH-monotonicity is still prominent in a microcanonical ensemble defined in an energy window $\delta E$ of the order $O(1/\sqrt{G}_{N})$.
@@
    22-d holographic CFT does not possess all the features of ETH-monotonicity. In particular, $\eta=3/2,\kappa=\pi\omega/4$ for $\Delta$ equal to integers and half integers. So, $f(\bar{E})=f_{0}e^{-2\kappa/\sqrt{\tilde{G}_{N}\bar{E}}}$ and the energy gains $\Delta E_{\beta}$ and $\Delta E_{n}$ vanish exponentially as $\bar{E}\to 0$. Moreover, quantum fluctuations numerically dominate ETH-monotonicity in $2$-d holographic CFT, unlike in higher-d CFTs.
@@
smaller black hole microstates in higher dimensions have stronger contribution from ETH-monotonicity to the second law of thermodynamics. In the smallest size limit, the black hole curvature singularity is constructed of microstates for which ETH-monotonicity starts competing with the entropic factor. In summary,
@@
Substituting these values in (3.12), we get (4.2). Surprisingly, it is independent of $\omega$ because $\kappa$ is independent of $\omega$ in these higher-d CFTs. $\overline{\Delta E}_{mc}$ is proportional to the square root of the Kretschmann scalar $\sqrt{K}=\sqrt{R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}}$ in $4$-D black holes. In higher-D black holes, $\overline{\Delta E}_{mc}$ measures higher order curvature invariants. The Kretschmann scalar at $r=r_{h}\to 0$ in different dimensions are
@@
It is worthwhile to note that if we naively take $r_{h}$ to the Planck length scale $l_{P}=G_{N}^{1/(d-1)}$, $\overline{\Delta E}_{mc}$ is of the order of one. So naively, the contributions to the gain in energy $\Delta E_{mc}$ from the entropic factor and the ETH-monotonicity become equal. One has to deal with quantum gravity at the Planck length scale. But we expect that ETH-monotonicity will persist even at this length scale, irrespective of the nature of quantum gravity. It is a property of many-body quantum chaotic system which becomes more prominent with decreasing system size, unlike other physical properties which are usually more predictable with increasing system size. It is generally agreed that the ultimate theory of quantum gravity will be a highly interacting chaotic theory. We have prominently observed ETH-monotonicity even in systems with 12 to 25 qubits ETHmono ; sorokhaibam2025ethmono2D .
In hindsight, our result is not surprising. Because canonical ensemble is ill-defined for small AdS black hole, as in flat spacetime. The entropy $S(\bar{E})$ grows super-linearly with increasing energy $\bar{E}$. So, the overarching idea is to exploit the inequivalence of ensembles to identify geometry with strong curvature. This is possible because of the surprising result that the extra gain in energy in microcanonical ensemble is always positive.
A small white hole is again a microstate with ETH-monotonicity dominance but at negative inverse temperature. A small perturbation will lead to the lost of an excess amount of energy of the system, more than what one expects from only entropic consideration.
@@
It is well known that certain geodesics can explore the black hole singularity Fidkowski:2003nf ; Festuccia:2005pi ; Rodriguez-Gomez:2021pfh ; Horowitz:2023ury ; Ceplak:2024bja ; Afkhami-Jeddi:2025wra ; Dodelson:2025jff . Other related works are Grinberg:2020fdj ; David:2022nfn ; Singhi:2024sdr ; Jia:2025jbi . But the exact nature of the singularity has been as elusive as ever. Our approach is new and radically different from the geodesic analysis. We are identifying microstates using the $f$-function in ETH statement.

### 25 Mar 2026 v4

@@ 3 ETH-monotonicity
$f(\bar{E})$ is closely related with the second law of thermodynamics. In ETHmono ; ETHmono2D , we found some interesting constraints on $\bar{E}$-dependence of the $f$-function in (1.1). These constraints are collectively called ETH-monotonicity. These constraints arise from a non-trivial relation between Berry’s conjecture and the second law of thermodynamics in the form of the Kelvin statement. It implies that, when the system is perturbed starting from an energy eigenstate or a microcanonical ensemble, more energy is absorbed than an equivalent canonical ensemble. The systems considered were lattice models with various energy scales. So, we will provide motivation and redefine ETH-monotonicity for holographic CFTs below. The main implication is that an energy eigenstate absorbs more energy than an equivalent canonical ensemble, when the system is perturbed. The low energy eigenstates are of particular interest. The high energy eigenstates gain energy same as their equivalent canonical ensembles due to equivalence of ensembles.
Now we will define ETH-monotonicity for holographic CFTs. It is a set of desirable criteria for the function $f(\bar{E})$. We will also provide the motivations for these criteria. ETH-monotonicity for holographic CFTs are the following criteria:
@@
(3.2) is the most severe constraint, $\eta=1$ is necessary. We emphasize that while taking $\bar{E}\to 0$ we are not shrinking the horizon radius to the string length scale $l_{s}$ or the Planck length scale $l_{P}$. In other word, the ground state $\bar{E}=0$ corresponds to the smallest black hole in the Hilbert space of string theory. We will briefly comment on the quantum gravity limit in the next section 4. $\bar{E}=0$ is the ground state because the Casimir energy does not participate in the dynamical set-up that we are considering. All CFTs do not have any intrinsic system size scale. So, R3 is formulated in terms of $\bar{E}$. R3 was originally formulated in terms of the degrees of freedom $L^{d}$, where $L$ is the linear system size and $d$ is the spatial dimension of the system.
For the Kelvin statement of the second law, the perturbation is by a source term $\lambda(t)\mathcal{O}$. Starting from a canonical ensemble with energy $\bar{E}_{\beta}$, the gain in energy by the system at leading order in perturbative expansion is ETHmono
@@
$\tilde{\lambda}(\omega)$ is the Fourier transform of $\lambda(t)$. For a hermitian operator, $\lambda(t)$ is real and $\tilde{\lambda}(\omega)=\tilde{\lambda}(-\omega)^{*}$. The above quantity is identically positive (negative or zero) for positive (negative or infinite) temperature. So, the second law of thermodynamics is guaranteed in canonical ensembles (in general, for any passive density matrix) Polkovnikov_2008 . The integral is finite because the $f$-function falls at least as fast as $e^{-|\beta\omega|/4}$, otherwise the theory is not UV complete. $\Delta E_{\beta}$ is of order $O(1)$ because we are perturbing the system with a light operator. Note that we have solved the backreacted geometry at leading order in perturbation theory. The gain in energy is the change in the mass of the black hole. We are performing the calculation from the CFT side. It is possible because of our assumption that holographic CFTs satisfy ETH.
@@
We have used Taylor expansions of $S(E_{n}+\omega)$ and $S(E_{n}+\omega/2)$ about the energy of the initial state $E_{n}$, $\omega=E_{m}-E_{n}$, and $\beta=S^{\prime}(\bar{E})|_{\bar{E}=E_{n}}$ is the effective inverse temperature at $E_{n}$. The difference with the expression of $\Delta E_{\beta}$ is in the $\bar{E}$-dependence. For a finite system (not thermodynamic limit), we will find that $f(\bar{E})$ is a monotonically increasing function of $\bar{E}$. It is also the criteria R2 that we set down in ETH-monotonicity. So the system gains more energy if the initial state is an eigenstate, over and above the entropic contribution. This extra gain in energy is due to ETH-monotonicity. We have neglected the contributions of the higher derivative terms $S^{\prime\prime}(\bar{E})$, $S^{\prime\prime\prime}(\bar{E})$, etc. They are suppressed by increasing positive powers of $G_{N}$ factor. But $S^{\prime\prime}(\bar{E})$ can compete with ETH-monotonicity. This term will be simply called quantum fluctuation. We will study it in details in section 5. These fluctuations, from both $f(E_{n}+\omega/2,\omega)$ and $S^{\prime\prime}(\bar{E})$, have been observed earlier (see around eqn (211) in DAlessio:2015qtq ). They were called “fluctuation within each eigenstate”. But the functional form of $f(\bar{E})$ has not been studied anywhere before ETHmono ; ETHmono2D . The “fluctuation from energy fluctuations” is the microcanonical fluctuations $\langle\mathcal{H}_{0}^{2}\rangle-\langle\mathcal{H}_{0}\rangle^{2}$ measurable in experiments. $\mathcal{H}_{0}$ is the Hamiltonian of the system. This fluctuation is zero for an energy eigenstate, but non-zero for a microcanonical ensemble. So either we work with an energy eigenstate or simply concentrate on expectation value of the energy when working with a microcanonical ensemble.
(3.7) also holds true in a microcanonical ensemble with energy expectation value $\bar{E}_{mc}$, see Appendix A for details. The microcanonical energy window $\delta$ is of the order of $O(1/\sqrt{G_{N}})$. But $\delta$ is usually taken to be parametrically much smaller than $\sqrt{G_{N}}\bar{E}_{mc}\sim 1/\sqrt{G_{N}}$ Srednicki_1999 . The correction from the microcanonical energy window $\delta$ arises from two different terms - one from the entropy exponentials and the other from $f(\bar{E})$. But both corrections are suppressed by $e^{-1/G_{N}}$. So $\delta$ does not affect the extra gain in energy from ETH-monotonicity in any dimension. This is expected because ETH-monotonicity is one of the manifestations of inequivalence of ensembles and we are working with microcanonical ensemble. The energy fluctuation $\langle\mathcal{H}_{0}^{2}\rangle-\langle\mathcal{H}_{0}\rangle^{2}\sim\delta\sim O(1/\sqrt{G_{N}})$ is significant for experimental measurements. As we will see, we are looking for $O(G^{1}_{N})$ effects which are very small compared to this energy fluctuation in measurements. So, one has to perform a large number of very high precision energy measurements to discern the $O(G^{1}_{N})$ effects (no free lunch). But we will concentrate on the expectation value and ignore the measurement fluctuations. In other words, our priority is to keep track of the minute $O(G^{1}_{N})$ effects theoretically. The microcanonical energy fluctuation is simply an experimental challenge.
In holographic CFTs, $\bar{E}$ is order $1/\tilde{G}_{N}$ times larger than $\omega$. So, we can Taylor expand $f(E_{n}+\omega/2,\omega)$ about $(E_{n},\omega)$ in (3.7). Using the Taylor expansion, we can extract the extra gain in energy from ETH-monotonicity when the system is initially in a microcanonical ensemble. Let us denote that quantity by $\Delta E_{mc,f}$ and it is given by
@@
The extra gain in energy is order $O(G^{1}_{N})$ because it is controlled by $f^{\prime}(\bar{E})/f(\bar{E})$, compared to $\Delta E_{\beta}$. So $f^{\prime}(\bar{E})/f(\bar{E})$ is the main object of our interest.
Pure states: In place of the microcanonical ensemble, one can also consider pure states $|\Psi\rangle=\sum_{n}c_{n}|n\rangle$ with $|n\rangle$ within the energy window $\delta$ of the order $O(1/\sqrt{G_{N}})$. These states have been called typical states in Srednicki_1999 . For the change in energy, the calculation is same as in the microcanonical ensemble because the cross terms $\mathcal{O}_{n_{1}m}\mathcal{O}_{mn_{2}}$ with $n_{1}\neq n_{2}$ in (3.5) drops out due to the pseudo random factor $R_{n_{1}m}R_{mn_{2}}$ averaging out to zero.
The motivations for ETH-monotonicity are as follows. R1 follows from equivalence of ensembles in the thermodynamic limit $\bar{E}\to\infty$. (3.3) of R3 is the mathematical expression. It implies $\Delta E_{mc}\to\Delta E_{\beta}$ as a power law in $\bar{E}_{mc}$ when taking the thermodynamic limit. $f(\bar{E})$ quantifies fluctuations (not relative fluctuations) at the energy $\bar{E}$. So, it is expected to be an increasing function of $\bar{E}$. This gives R2. Many observables in chaotic lattice models have been found to obey R2 in ETHmono ; ETHmono2D .
We have special interest in the limit $\bar{E}\to 0$ because it corresponds to small black holes with high curvature at the horizon, and ultimately the black hole singularity. Naively, we expect that the extra gain in energy would persist in the limit $\bar{E}\to 0$ if $\eta>0$ in (3.2). $f^{\prime}/f$ will diverge as $\kappa\bar{E}^{-\eta}$. It ensures that $\Delta E_{mc,f}$ increases compared to $\Delta E_{\beta}$ (or $\Delta E_{mc,f}/\Delta E_{\beta}$ increases). For example, we will find that 2d2-d holographic CFT has $\eta=3/2,\kappa=\pi\omega/4$. But with this value of $\eta$, $f(\bar{E})$ is
@@
where $f_{0}>0$ is the integration constant. The problem with this function is that $\Delta E_{mc},\Delta E_{\beta}\sim e^{-1/\bar{E}}$ vanishes exponentially fast. There could be power-law dependent factor of $\bar{E}$ but it is not important for $f^{\prime}(\bar{E})/f(\bar{E})$ in the presence of the above exponential. On the other hand, if $\eta=1$, then $f(\bar{E})$ is
$f(\bar{E})=f_{0}\bar{E}^{\kappa}.$ (3.11)
So, $f(\bar{E})$ vanishes as a power law in $\bar{E}$. Another useful consequence is that $\kappa$ is generally expected to be independent of $\omega$, on general consideration. In the present physics problem, we do not expect a function like $\bar{E}^{\beta\omega}$ or $\bar{E}^{\log(\beta\omega)}$ where $\bar{E}$ is the energy of the system and $\omega$ is the frequency of the perturbation or linear response. This remark on $\kappa$ will have an interesting consequence later on.
If $\eta$ is less than $1$, say $\eta=1/2$, then $f(\bar{E})=f_{0}e^{2\kappa\sqrt{\tilde{G}_{N}\bar{E}}}$ which is not possible because $\bar{E}=0$ is the ground state. $f(\bar{E})$ signifies the strength of quantum fluctuations mimicking thermal fluctuations Srednicki_1999 and it is expected to vanish for the ground state. So in conclusion, $\eta=1$ is a very strict constraint in (3.2).
Note that $\eta>0$, in particular $\eta=1$, implies that ETH-monotonicity can start competing with the entropic contribution in the limit $\bar{E}\to 0$. We can examine this competition by calculating the relative extra gain in energy at a particular frequency $\omega$ defined by
@@ 4 Summary of results
We use the $\bar{E}$-dependence of $f(\bar{E},\omega)$ in the ETH statement (1.1) to identify and examine black hole microstates. In particular, we are interested in the microstates of small black holes which have high curvature at the horizon. We set down a set of criteria on the $\bar{E}$-dependence which we called ETH-monotonicity. Our technical results are:
@@
    ETH-monotonicity is still prominent in a microcanonical ensemble defined in an energy window $\delta$ of the order $O(1/\sqrt{G}_{N})$. In Higher-d holographic CFTs, the relative extra gain in energy $\overline{\Delta E}_{mc}$ measures the curvature at the black hole horizon of small black holes,
    $\lim_{\bar{E}\to 0}\overline{\Delta E}_{mc}=\frac{\kappa(d-2)\,\tilde{G}_{N}}{2\pi r_{h}^{d-1}}=\frac{\kappa(D-3)\,\tilde{G}_{N}}{2\pi r_{h}^{D-2}},\quad\text{with}\quad D=d+1\geq 4.$ (4.2)
    ETH-monotonicity also dominates the quantum fluctuation in the small black hole microstates.
@@
    22-d holographic CFT does not possess all the features of ETH-monotonicity. In particular, $\eta=3/2,\kappa=\pi\omega/4$ for $\Delta$ equal to integers and half integers. So, $f(\bar{E})=f_{0}e^{-2\kappa/\sqrt{\tilde{G}_{N}\bar{E}}}$ and the energy gains $\Delta E_{\beta}$ and $\Delta E_{mc}$ vanish exponentially as $\bar{E}\to 0$. Moreover, the quantum fluctuation negates and numerically dominates ETH-monotonicity in $2$-d holographic CFT.
@@
smaller black hole microstates in higher dimensions have stronger contribution from ETH-monotonicity to the second law of thermodynamics. In the smallest size limit, the black hole curvature singularity is constructed of microstates for which ETH-monotonicity starts competing with the entropic contribution. In summary,
@@
Substituting these values in (3.12), we get (4.2). The negative specific heat of small black holes implies that $\omega\coth(\beta\omega/2)\sim 2/\beta$ in (3.12) as $\beta\to 0$. $\kappa$ is also independent of $\omega$ in these higher-d CFTs. So the expression of $\overline{\Delta E}_{mc}$ is fixed by dimensional analysis after the $\omega$-dependence dropped out. As we have remarked above, we expect that $\kappa$ is independent of $\omega$ if $\eta=1$, or in general, if ETH-monotonicity is satisfied. We expect that this result would generalize to other black hole geometries with negative specific heat and satisfying ETH-monotonicity.
$\overline{\Delta E}_{mc}$ is proportional to the square root of the Kretschmann scalar $\sqrt{K}=\sqrt{R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}}$ in $4$-D black holes. In higher-D black holes, $\overline{\Delta E}_{mc}$ measures higher order curvature invariants. The Kretschmann scalar at $r=r_{h}\to 0$ in different dimensions are
@@
It is worthwhile to note that if we naively take $r_{h}$ to the Planck length scale $l_{P}=G_{N}^{1/(d-1)}$, $\overline{\Delta E}_{mc}$ is of the order of one. So naively, the entropic contribution $\Delta E_{\beta}$ and the contribution from ETH-monotonicity become equal. One has to deal with quantum gravity at the Planck length scale. But we expect that ETH-monotonicity will persist even at this length scale, irrespective of the nature of quantum gravity. It is a property of many-body quantum chaotic system which becomes more prominent with decreasing system size, unlike other physical properties which are usually more predictable with increasing system size. It is generally agreed that the ultimate theory of quantum gravity will be a highly interacting chaotic theory. We have prominently observed ETH-monotonicity even in systems with 12 to 25 qubits ETHmono ; ETHmono2D .
In hindsight, our result is not surprising. Because canonical ensemble is ill-defined for small AdS black holes, as for black holes in flat spacetime. The entropy $S(\bar{E})$ grows super-linearly with increasing energy $\bar{E}$. So, one has to work with microcanonical ensemble. Then, the overarching idea is to exploit the inequivalence of ensembles to identify geometry with strong curvature. This is possible because of the surprising result that the extra gain in energy in microcanonical ensemble is always positive. We emphasize that our results are neither conjectures nor proposals. They follow from precise and concrete calculations. We have analysed all competing terms at order $O(G^{1}_{N})$.
A large “white hole” [^2] We do not mean the white hole in the Penrose diagram of an eternal black hole. We mean the microstates in the upper half of the finite energy spectrum of quantum gravity. The opposite of small black holes are not small “white holes”. They are the highest energy microstates with small entropy, so large “white holes”. is again a microstate with ETH-monotonicity dominance but at negative inverse temperature. A small perturbation will lead to the lost of an excess amount of energy of the system, more than what one expects from only entropic consideration. The entropy will decrease with increasing energy for such “white holes”. So, $f(\bar{E})$ will be a decreasing function of $\bar{E}$ in this region of the energy spectrum.